OpenStudy (katherinesmith):

How many ways are there to choose a committee of 7 people from a group of 10 people? Please help, medal will be awarded

OpenStudy (wolf1728):

The formula for choosing 'k' people from a larger group of 'n' people: Combinations = n!/(k!)*(n-k)! Combinations = 10! / 7! * (3)! Combinations = 10 * 9 * 8 * 7! / 3! * 7! Combinations = 10 * 9 * 8 / 3! Combinations = 720 / 6 Combinations = 120

OpenStudy (katherinesmith):

thank you! can you help me with another similar one

OpenStudy (wolf1728):

Well okay

OpenStudy (katherinesmith):

Find the number of permutations in the word arithmetic.

OpenStudy (wolf1728):

arithmetic is 10 letters duplicate letters are 'i' 't' Combinations = 10! / (2! * 2!) Combinations = 10 * 9 * 8 * 7 * 6 * 5 * 2 * 3 Combinations = 907,200